OpenStudy (anonymous):

7 - 7(x - 7) > -4 + 5x

OpenStudy (lgbasallote):

7 - 7x + 49 > -4 + 5x 56 > -4 + 12x 60 > 12x

OpenStudy (anonymous):

=7-7x +7 > -4+5x =14+4>5x+7x =18>12x =x<18/12 =x<3/2

OpenStudy (anonymous):

7-7x+49>-4+5x 56+4>5x+7x 60>12x so....x<60/12 x<5